At the lake Curtis caught three fish .He wanted to wiegh the fish individually but his scale was broken -- it could only read weights between 5 and 10 pounds So he weighed the large and medium fish togethter and got 7.8 pounds The large and small fish weighed 7.0 pounds and the small medium fish were 5.6 pounds. How much did each fish weigh individually? What was the total weight of all three fish?
Answer by stanbon(57222)

At the lake Curtis caught three fish .He wanted to wiegh the fish individually but his scale was broken -- it could only read weights between 5 and 10 pounds So he weighed the large and medium fish together and got 7.8 pounds The large and small fish weighed 7.0 pounds and the small medium fish were 5.6 pounds. How much did each fish weigh individually? What was the total weight of all three fish?
-----------
Let the three fish be L, M, and S in weight.
----------------------------------
EQUATION:
L + M + 0 = 7.8
L + 0 + S = 7.0
0 + M + S = 5.6
-------------------
Solve by any method to get:
L = 4.6 lbs.
M = 3.2 lbs.
S = 2.4 lbs.
=====================
Cheers,
Stan H.